OpenStudy (moongazer):

The speed of a car is 80 miles per hour. Find the radius of the car wheels in inches if the car makes 300 rev/min. I think it has something to do with linear and angular speed :)

OpenStudy (moongazer):

please help :)

OpenStudy (saifoo.khan):

Your first target is to find it's circumference.

OpenStudy (anonymous):

Calculate the speed per minute by dividing the wheel speed Multiply the miles per minute speed times 60 to convert the speed to miles per hour (mph).

OpenStudy (anonymous):

Thats all!

OpenStudy (moongazer):

how do I do that?

OpenStudy (moongazer):

@saifoo.khan What should I do next?

OpenStudy (saifoo.khan):

i will brb.

OpenStudy (moongazer):

ok :)

OpenStudy (anonymous):

Moon the tips are given still you are unable to solve!

OpenStudy (ganpat):

yes you r right.. V = rw where V = linear velocity and w = angular velocity...

OpenStudy (anonymous):

See they did it! especially @Ganpat

OpenStudy (moongazer):

I'll try to do it :)

OpenStudy (anonymous):

Good Luck

OpenStudy (ganpat):

v = 80 miles/ hr r = find ?? w = (2 * pi * N )/ 60.. N = 300 rev /min = 300 * 60 = 1800 rev / hr.. Now put everything in the formula and look for the unknown.. i.e. r.. V = rw ...

OpenStudy (moongazer):

@Ganpat I did a slightly different method. I got r=140.8 inches. is it right?

OpenStudy (moongazer):

I rechecked it I got r=44.82 inches

OpenStudy (anonymous):

yes it is correct!

OpenStudy (moongazer):

which one?

OpenStudy (anonymous):

44.82 inches

OpenStudy (moongazer):

Thanks :)

OpenStudy (moongazer):

I forgot the pi with my first answer :)